travis is seven years younger than his sister if you take double the sum of both their ages they are 36
Answer by ankor@dixie-net.com(22740)

travis is seven years younger than his sister
t = s - 7
if you take double the sum of both their ages, they are 36
2(t + s) = 36
simplify, divide by 2
t + s = 18
replace t with (s-7)
s - 7 + s = 18
s + s = 18 + 7
2s = 25
s = 25/2
s = 12.5 is the sisters age
then
5.5 = t's age
;
:
See if that checks out.
2(5.5 + 12.5) = 36